# RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model

### v1

@@ Abstract
Human Action Recognition (HAR) encompasses the task of monitoring human activities across various domains, including but not limited to medical, educational, entertainment, visual surveillance, video retrieval, and the identification of anomalous activities. Over the past decade, the field of HAR has witnessed substantial progress by leveraging Convolutional Neural Networks (CNNs) to effectively extract and comprehend intricate information, thereby enhancing the overall performance of HAR systems. Recently, the domain of computer vision has witnessed the emergence of Vision Transformers (ViTs) as a potent solution. The efficacy of transformer architecture has been validated beyond the confines of image analysis, extending their applicability to diverse video-related tasks. Notably, within this landscape, the research community has shown keen interest in HAR, acknowledging its manifold utility and widespread adoption across various domains. This article aims to present an encompassing survey that focuses on CNNs and the evolution of Recurrent Neural Networks (RNNs) to ViTs given their importance in the domain of HAR. By conducting a thorough examination of existing literature and exploring emerging trends, this study undertakes a critical analysis and synthesis of the accumulated knowledge in this field. Additionally, it investigates the ongoing efforts to develop hybrid approaches. Following this direction, this article presents a novel hybrid model that seeks to integrate the inherent strengths of CNNs and ViTs.
@@
Human action recognition; convolutional neural networks; recurrent neural networks; vision transformers; deep learning; video classification
@@ 1 Introduction
The advent of Transformer[5] architecture has sparked a paradigm shift in deep learning, leading researchers to swiftly adopt and adapt these architectures to enhance the accuracy and efficiency of various computer vision tasks including action recognition. Consequently, there is a compelling need for a recent review that comprehensively examines the state-of-the-art research including Vision Transformers (ViTs) [6], a special Transformer for computer vision tasks, and Hybrid models in addition to CNNs. Such a review would serve as a crucial guiding resource to shape the future research directions with Transformer and CNN-Transformer hybrid architectures beside CNNs which previously were seen as unique and influential models for action recognition.
@@
By employing a multi-head self-attention layer, the Transformer model computes sequence representations by effectively aligning words within the sequence with other words in the same sequence [1]. This approach outperforms traditional convolutional and recursive operations in terms of representation quality while utilizing fewer computational resources. As a consequence, the Transformer architecture diverges from conventional convolutional and recursive methods, favoring a more focused utilization of multiple processing nodes. The incorporation of multi-head attention allows the Transformer model to collectively learn a range of representations from diverse perspectives through the collaboration of multiple attention layers. Inspired by Transformers, several Natural Language Processing (NLP) tasks have achieved remarkable performance, reaching human-level capabilities, as exemplified by models such as GPT [7] and BERT [8].
@@
The remarkable achievements of Transformers in handling sequential data, particularly in the domain of NLP, have prompted the exploration and advancement of ViTs [6]. These novel models have demonstrated comparable or even superior performance compared to CNNs in the context of image recognition tasks, especially when operating on vast datasets such as Imagenet [9, 10, 11]. This observation signifies a noteworthy shift in the field, wherein ViTs possess the potential to supplant the established dominance of CNNs in computer vision, mirroring the displacement witnessed in the case of Recurrent Neural Networks (RNNs) [1]. The achievements of Transformer models have engendered considerable scholarly interest within the computer vision research community, prompting rigorous exploration of their efficacy in pure computer vision tasks.
@@
The natural progression in the advancement of ViTs has led to the logical exploration of video recognition tasks. Unlike image recognition, video recognition focuses on the complex challenge of identifying and understanding events within video sequences, including the recognition of human actions. The main contributions of this paper is as follows.
@@
  • We present a thorough review of the CNNs, RNNs, and ViTs. This review examines the evolution from traditional methods to the latest advancements in neural network architectures.
  • We present an extensive examination of existing literature related to action recognition.
@@
  • We propose a novel hybrid model integrating the strengths of CNNs and ViTs. In addition, we provide a detailed performance comparison the proposed hybrid model against existing models. The analysis highlights the model’s efficacy in handling complex HAR tasks with improved accuracy and efficiency.
@@
  • We also discuss emerging trends and the future direction of HAR technologies, emphasizing the importance of hybrid models in enhancing the interpretability and robustness of action recognition systems.
@@
The paper is structured as follows: Section 2 delves into the background, covering foundational concepts and technologies crucial to HAR, including CNNs, RNNs, and ViTs. Section 3 reviews related work, highlighting the chronological evolution of HAR deep learning technologies. The novel hybrid model combining CNNs and ViTs is introduced in Section 4, followed by Section 5, which details the experimental setup and evaluation of the model. Section 6 discusses the challenges and their implications for future directions. Finally, Section 7 concludes the paper.
@@ 2.1 From Vanilla RNN to Attention-based Transformers
The fundamental aim of the Transformers lies in effectively addressing time series data tasks, placing a distinct emphasis on NLP as one of its primary objectives. To address these tasks, RNNs have emerged as a prominent deep learning technique. However, the development of the Transformers did not occur in isolation; rather, it was the result of a series of iterative optimisations that were progressively integrated into the existing RNN structures. This section aims to elucidate the underlying RNN techniques and the subsequent advancements that were incorporated into them, ultimately leading to the final formulation of the Transformer architecture.
Notation. Here, we establish common notations for RNN architectures including Vanilla RNNs, Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) to streamline discussions in subsequent sections. In these architectures, each iteration involves a cell that sequentially processes an input embedding $\boldsymbol{x}_{t}\in\mathbb{R}^{n_{x}}$ and retains information from the previous sequence through the hidden state $\boldsymbol{h}_{t-1}\in\mathbb{R}^{n_{h}}$ using weight matrices $\boldsymbol{W}\in\mathbb{R}^{n_{h}\times n_{h}}$ and $\boldsymbol{U}\in\mathbb{R}^{n_{h}\times n_{x}}$. The $\boldsymbol{W}$-like matrices encompasses all weights related to hidden-to-hidden connections, while $\boldsymbol{U}$-like matrices encompasses all weight matrices related to input-to-hidden connections. Additionally, bias terms are represented by $\boldsymbol{b}$-like vectors. Each cell produces a new hidden state $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ as its output.
@@
Vanilla RNNs effectively incorporate short-term dependencies of temporal order and past inputs in a meaningful manner. However, they are characterised by certain limitations. Firstly, due to their intrinsic sequential nature, RNNs pose challenges in parallelizing computations [14]. Consequently, this limitation can impose restrictions on the overall speed and scalability of the network. Secondly, when processing lengthy sequences, the issue of exploding or vanishing gradients may arise, thereby impeding the stable training of the network [15].
@@
LSTM Hochreiter and Schmidhuber (1997) introduced the Long Short-Term Memory (LSTM) cell as a solution to address the issue of long-term dependencies and to mitigate the challenge of interdependencies among successive steps [16]. LSTM architecture incorporates a distinct component known as the cell state $\boldsymbol{c}_{t}\in\mathbb{R}^{n_{h}}$, illustrated in Figure 1(b). Analogous to a freeway, this cell state facilitates the smooth flow of information, ensuring that it can readily traverse without undergoing significant alterations.
In 2000, Gers, Schmidhuber, and Cummins [17] made modifications to the initial LSTM architecture by incorporating a forget gate within the cell structure. The mathematical expressions describing this modified LSTM cell are derived from its inner connections. Hence, the LSTM cell can be formally represented based on the depicted interconnections as follows:
@@
    where $\boldsymbol{f}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the forget gate and $\sigma$ is sigmoid activation function.
@@
    where $\boldsymbol{i}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the sigmoid activation function and
    $\boldsymbol{\tilde{c}}_{t}=\tanh(\boldsymbol{W}_{\tilde{c}}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{\tilde{c}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{c}})$ (4)
@@
    $\boldsymbol{c}_{t}=\boldsymbol{f}_{t}\odot\boldsymbol{c}_{t-1}+\boldsymbol{i}_{t}\odot\boldsymbol{\tilde{c}}_{t}$ (5)
    where $\boldsymbol{c}_{t-1}\in\mathbb{R}^{n_{h}}$ is the previous cell state.
@@
    $\boldsymbol{o_{t}}=\sigma(\boldsymbol{W}_{o}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{o}\boldsymbol{x}_{t}+\boldsymbol{b}_{o})$ (6)
    where $\boldsymbol{o}_{t}\in\mathbb{R}^{n_{h}}$ is the output of output gate. Finally the updated hidden state,
@@
GRU The LSTM cell surpasses the learning capability of the conventional recurrent cell, yet the additional number of parameters escalates the computational load. Consequently, to address this concern, Chung et al. [18] introduced the gated recurrent unit (GRU), see Figure 2. GRU demonstrates comparable performance to LSTM while offering a more computationally efficient design with fewer weights. This is achieved by merging the cell state and the hidden state into “reset state” resulting in a simplified architecture. Furthermore, GRU combines the forget and input gates into an “update gate”, contributing to a more streamlined computational process. For further elaboration, GRU cell incorporates two essential gates. The first gate is the reset gate, which examines the previous hidden state $\boldsymbol{h}_{t-1}$ and the current input $\boldsymbol{x}_{t}$. It generates a vector $\boldsymbol{r}_{t}$ containing values between 00 and $1$, determining the extent to which past information in $\boldsymbol{h}_{t-1}$ should be disregarded. The second gate is the update gate, which governs the selection of information to either retain or discard when updating the new hidden state $\boldsymbol{h}_{t}$, based on the value of $\boldsymbol{r}_{t}$.
@@
    $\boldsymbol{z}_{t}=\sigma(\boldsymbol{W}_{z}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{z}\boldsymbol{x}_{t}+\boldsymbol{b}_{z})$ (8)
    where $\boldsymbol{z}_{t}\in\mathbb{R}^{n_{h}}$ is the output of update gate. A candidate activation for the subsequent step
@@
    where $\boldsymbol{\tilde{h}}_{t}\in\mathbb{R}^{n_{h}}$ and
    $\boldsymbol{r}_{t}=\sigma(\boldsymbol{W}_{r}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{r}\boldsymbol{x}_{t}+\boldsymbol{b}_{r})$ (10)
    where $\boldsymbol{r}_{t}\in\mathbb{R}^{n_{h}}$ is the output of reset gate.
  • Final Activation is a blend of the previous hidden state and the candidate activation, weighted by the update gate. This mechanism allows the GRU to effectively retain or replace old information with new information.
@@
    where $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ is the updated hidden state.
Types of RNNs RNNs were created with an internal memory mechanism that allows them to store and use information from previous outputs. This unique trait enables RNNs to retain important contextual information over time, enabling reasoned decision-making based on past results. There are four types of popular RNN variants that each serve different purposes across a variety of applications, see Figure 3.
The One-to-one is considered the simplest form of RNNs, where a single input corresponds to a single output. It operates with fixed input and output sizes, functioning similarly to a standard neural network.

### v2

# CNNs, RNNs and Transformers in Human Action Recognition: A Survey and a Hybrid Model

@@ Abstract
Human action recognition (HAR) encompasses the task of monitoring human activities across various domains, including but not limited to medical, educational, entertainment, visual surveillance, video retrieval, and the identification of anomalous activities. Over the past decade, the field of HAR has witnessed substantial progress by leveraging convolutional neural networks (CNNs) and recurrent neural networks (RNNs) to effectively extract and comprehend intricate information, thereby enhancing the overall performance of HAR systems. Recently, the domain of computer vision has witnessed the emergence of Vision Transformers (ViTs) as a potent solution. The efficacy of Transformer architecture has been validated beyond the confines of image analysis, extending their applicability to diverse video-related tasks. Notably, within this landscape, the research community has shown keen interest in HAR, acknowledging its manifold utility and widespread adoption across various domains. This article aims to present an encompassing survey that focuses on CNNs and the evolution of RNNs to ViTs given their importance in the domain of HAR. By conducting a thorough examination of existing literature and exploring emerging trends, this study undertakes a critical analysis and synthesis of the accumulated knowledge in this field. Additionally, it investigates the ongoing efforts to develop hybrid approaches. Following this direction, this article presents a novel hybrid model that seeks to integrate the inherent strengths of CNNs and ViTs.
@@
Human action recognition · Convolutional neural networks · Recurrent neural networks · Vision Transformers · Deep learning · Video classification
@@ 1 Introduction
The advent of Transformer architecture [96] has sparked a paradigm shift in deep learning. By employing a multi-head self-attention layer, the Transformer model computes sequence representations by effectively aligning words within the sequence with other words in the same sequence [93]. This approach outperforms traditional convolutional and recursive operations in terms of representation quality while utilizing fewer computational resources. As a consequence, the Transformer architecture diverges from conventional convolutional and recursive methods, favoring a more focused utilization of multiple processing nodes. The incorporation of multi-head attention allows the Transformer model to collectively learn a range of representations from diverse perspectives through the collaboration of multiple attention layers. Inspired by Transformers, many natural language processing (NLP) tasks have achieved remarkable performance, reaching human-level capabilities, as exemplified by models such as GPT [12] and BERT [24].
@@
The remarkable achievements of Transformers in handling sequential data, particularly in the domain of NLP, have prompted the exploration and advancement of Vision Transformer (ViT) [29] (a special Transformer for computer vision tasks). ViTs have demonstrated comparable or even superior performance compared to CNNs in the context of image recognition tasks, especially when operating on vast datasets such as ImageNet [41, 66, 54]. This observation signifies a noteworthy shift in the field, wherein ViTs possess the potential to supplant the established dominance of CNNs in computer vision, mirroring the displacement witnessed in the case of recurrent neural networks (RNNs) [93]. The achievements of Transformer models have engendered considerable scholarly interest within the computer vision research community, prompting rigorous exploration of their efficacy in pure computer vision tasks.
@@
The natural progression in the advancement of ViTs has led to the logical exploration of video recognition tasks. Unlike image recognition, video recognition focuses on the complex challenge of identifying and understanding events within video sequences, including the recognition of human actions. Consequently, there is a compelling need for a recent review that comprehensively examines the state-of-the-art research including ViTs and hybrid models in addition to CNNs and RNNs for HAR. Such a review would serve as a crucial guiding resource to shape the future research directions with Transformer and CNN-Transformer hybrid architectures beside CNNs which previously were seen as unique and influential models for HAR. The main contributions of this paper is as follows.
@@
  • We present a thorough review of the CNNs, RNNs and ViTs. This review examines the evolution from traditional methods to the latest advancements in neural network architectures.
  • We present an extensive examination of existing literature related to HAR.
@@
  • We propose a novel hybrid model integrating the strengths of CNNs and ViTs. In addition, we provide a detailed performance comparison of the proposed hybrid model against existing models. The analysis highlights the model’s efficacy in handling complex HAR tasks with improved accuracy and efficiency.
@@
  • We also discuss emerging trends and the future direction of HAR technologies, emphasizing the importance of hybrid models in enhancing the interpretability and robustness of HAR systems.
@@
The paper is structured as follows. Section 2 delves into the background, covering foundational concepts and technologies crucial to HAR, including CNNs, RNNs and ViTs, highlighting the chronological evolution of HAR deep learning technologies. Section 3 thoroughly reviews related HAR works with a brief discussion. A novel hybrid model combining CNNs and ViTs is proposed in Section 4, including the details of the experimental setup and the results. Section 5 discusses the challenges and their implications for future directions in HAR. Finally, Section 6 concludes the paper.
@@ 2 Background
This section provides a chronological and technical overview of three fundamental types of neural networks: CNNs, RNNs, and Transformers. CNNs, introduced in the late 1980s, revolutionized image processing by leveraging local connectivity and shared weights to efficiently detect spatial hierarchies in data. As the field progressed, RNNs emerged in the 1990s, addressing the need for modeling sequential data through their ability to maintain temporal dependencies across sequences. The advent of Transformers in 2017 marked a paradigm shift by utilizing self-attention mechanisms to capture global relationships in data more effectively, thereby enhancing performance in a wide array of tasks beyond sequential data. This background section will delve into the technical intricacies and evolutionary trajectory of these architectures, highlighting their contributions and transitions in the realm of deep learning.
2.1 CNNs
The evolution of CNNs has been remarkable since their introduction in the 1980s. Originally, CNNs were designed to process static images, primarily focusing on spatial recognition tasks such as object and pattern recognition. The initial idea was to build layers of convolutional filters that would apply various operations to the image to extract features like edges, textures, and shapes. This structure proved highly effective for tasks like image classification, object detection, image segmentation and more in computer vision.
The Neocognitron [34], developed by Kunihiko Fukushima, presented an early example of neural networks incorporating convolutional operations for image processing, setting the foundations for subsequent progress. Shortly after, Yann LeCun and collaborators introduced LeNet-5 [61], a key architecture designed for handwritten digit recognition, showcasing the effectiveness of convolutional layers in pattern recognition tasks. The progress of CNNs reached a turning point in the mid-2010s with the introduction of models like AlexNet [58], showcasing their potential in image classification tasks. Alongside architectural innovations, this milestone was achieved thanks to access to large datasets, notably, ImageNet [23], and computational improvements, including the rise of graphics processing units (GPUs) for parallel computing. Large-scale datasets provided the diversity and complexity necessary for training deep networks, while enhanced computational power accelerated the training of sophisticated CNN architectures.
2.1.1 Spatio-Temporal CNNs
As CNNs excelled in spatial tasks, researchers began exploring their potential in handling temporal data, such as video and time-series analysis. The challenge was to incorporate the dimension of time into the inherently spatial architecture of CNNs. To address this task, spatio-temporal CNNs were developed. These networks extend traditional CNN architectures by adding a temporal component to analyze dynamic behaviors across time frames. Several approaches have been utilized and main types are as follows.
3D convolution involves extending the 2D kernels to 3D, allowing the network to perform convolution across both spatial and temporal dimensions. This approach is directly applied to video data where the third dimension represents time [42, 90]. The two-stream CNNs involve running two parallel CNN streams: one for spatial processing of individual frames and another for temporal processing, usually of optical flow, which captures motion between frames [81, 33]. RNNs with CNNs aim to combine CNNs for spatial processing with RNNs like long short-term memory (LSTM) or gated recurrent unit (GRU) to handle temporal dependencies. This hybrid model leverages CNNs’ ability to extract spatial features and RNNs’ capacity to manage temporal sequences effectively [111, 28].
@@ 2.2 From Vanilla RNN to Attention-Based Transformers
This section explores the evolution from RNNs to the Transformers, highlighting the progression in handling time series and sequence data. Initially, RNNs were the go-to deep learning technique for managing temporal tasks, effectively capturing sequential dependencies. However, the development of Transformers marked a significant leap forward, driven by a series of iterative improvements and optimizations that built upon the limitations of RNNs. Transformers, with their focus on NLP, introduced a novel attention mechanism that allows for more efficient and scalable processing of sequential data. By examining the foundational RNN techniques and the subsequent enhancements leading to the Transformer architecture, this section elucidates the transformative journey from traditional RNN models to the sophisticated attention-based frameworks that now dominate the field.
We firstly establish common notations for RNN architectures including vanilla RNNs, LSTM and GRU to streamline discussions in subsequent sections. In these architectures, each iteration involves a cell that sequentially processes an input embedding $\boldsymbol{x}_{t}\in\mathbb{R}^{n_{x}}$ and retains information from the previous sequence through the hidden state $\boldsymbol{h}_{t-1}\in\mathbb{R}^{n_{h}}$ using weight matrices $\boldsymbol{W}\in\mathbb{R}^{n_{h}\times n_{h}}$ and $\boldsymbol{U}\in\mathbb{R}^{n_{h}\times n_{x}}$. The $\boldsymbol{W}$-like matrices encompass all weights related to hidden-to-hidden connections, while $\boldsymbol{U}$-like matrices encompass all weight matrices related to input-to-hidden connections. Additionally, bias terms are represented by $\boldsymbol{b}$-like vectors. Each cell produces a new hidden state $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ as its output.
2.2.1 Vanilla RNNs
@@
Vanilla RNNs effectively incorporate short-term dependencies of temporal order and past inputs in a meaningful manner. However, they are characterized by certain limitations. Firstly, due to their intrinsic sequential nature, RNNs pose challenges in parallelized computations [40]. Consequently, this limitation can impose restrictions on the overall speed and scalability of the network. Secondly, when processing lengthy sequences, the issue of exploding or vanishing gradients may arise, thereby impeding the stable training of the network [9].
@@
2.2.2 LSTM
Hochreiter and Schmidhuber (1997) introduced the LSTM cell as a solution to address the issue of long-term dependencies and to mitigate the challenge of interdependencies among successive steps [44]. LSTM architecture incorporates a distinct component known as the cell state $\boldsymbol{c}_{t}\in\mathbb{R}^{n_{h}}$, illustrated in Figure 1(b). Analogous to a freeway, this cell state facilitates the smooth flow of information, ensuring that it can readily traverse without undergoing significant alterations.
[36] made modifications to the initial LSTM architecture by incorporating a forget gate within the cell structure. The mathematical expressions describing this modified LSTM cell are derived from its inner connections. Hence, the LSTM cell can be formally represented based on the depicted interconnections as follows.
@@ 2.2.2 LSTM
    where $\boldsymbol{f}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the forget gate and $\sigma$ is the sigmoid activation function.
@@
    where $\boldsymbol{i}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the sigmoid activation function; and
    $\boldsymbol{\tilde{c}}_{t}=\tanh(\boldsymbol{W}_{\tilde{c}}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{\tilde{c}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{c}}),$ (4)
@@
    $\boldsymbol{c}_{t}=\boldsymbol{f}_{t}\odot\boldsymbol{c}_{t-1}+\boldsymbol{i}_{t}\odot\boldsymbol{\tilde{c}}_{t},$ (5)
    where $\boldsymbol{c}_{t-1}\in\mathbb{R}^{n_{h}}$ is the previous cell state and $\odot$ is the Hadamard operator.
@@
    $\boldsymbol{o_{t}}=\sigma(\boldsymbol{W}_{o}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{o}\boldsymbol{x}_{t}+\boldsymbol{b}_{o}),$ (6)
    where $\boldsymbol{o}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the output gate. Finally the updated hidden state,
@@
2.2.3 GRU
The LSTM cell surpasses the learning capability of the conventional recurrent cell, yet the additional number of parameters escalates the computational load. Consequently, to address this concern, [18] introduced the GRU, see Figure 2. GRU demonstrates comparable performance to LSTM while offering a more computationally efficient design with fewer weights. This is achieved by merging the cell state and the hidden state into “reset state” resulting in a simplified architecture. Furthermore, GRU combines the forget and input gates into an “update gate”, contributing to a more streamlined computational process. For further elaboration, GRU cell incorporates two essential gates. The first gate is the reset gate, which examines the previous hidden state $\boldsymbol{h}_{t-1}$ and the current input $\boldsymbol{x}_{t}$. It generates a vector $\boldsymbol{r}_{t}$ containing values between 00 and $1$, determining the extent to which past information in $\boldsymbol{h}_{t-1}$ should be disregarded. The second gate is the update gate, which governs the selection of information to either retain or discard when updating the new hidden state $\boldsymbol{h}_{t}$, based on the value of $\boldsymbol{r}_{t}$.
@@ 2.2.3 GRU
    $\boldsymbol{z}_{t}=\sigma(\boldsymbol{W}_{z}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{z}\boldsymbol{x}_{t}+\boldsymbol{b}_{z}),$ (8)
    where $\boldsymbol{z}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the update gate. The output of the reset gate $\boldsymbol{r}_{t}\in\mathbb{R}^{n_{h}}$ is obtained by
    $\boldsymbol{r}_{t}=\sigma(\boldsymbol{W}_{r}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{r}\boldsymbol{x}_{t}+\boldsymbol{b}_{r}).$ (9)
    A candidate activation for the subsequent step is
@@
    where $\boldsymbol{\tilde{h}}_{t}\in\mathbb{R}^{n_{h}}$.
  • The final activation is a blend of the previous hidden state and the candidate activation, weighted by the update gate, i.e.,
@@
    where $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ is the updated hidden state. This mechanism allows the GRU to effectively retain or replace old information with new information.
@@
2.2.4 Types of RNNs
RNNs were created with an internal memory mechanism that allows them to store and use information from previous outputs. This unique trait enables RNNs to retain important contextual information over time, enabling reasoned decision-making based on past results. There are four types of popular RNN variants that each serve different purposes across a variety of applications, see Figure 3.